OpenStudy (azureilai):

can someone please show me how to convert the trigonometric form of 5(cos(0)+isin(0) into the standard form a+bi

OpenStudy (anonymous):

cos(0)=1 sin(0)=0

OpenStudy (azureilai):

I know the unit circle and the value of the cos and sin. My question is how to convert the trigonometric form into the standard a+bi form

OpenStudy (anonymous):

substitute in 1 for cos(0) and 0 for sin(0) 5(cos(0)+isin(0)) = 5(1+0i) = 5+0i

OpenStudy (azureilai):

Oh ok. I see, thanks. I didn't know you could substitute.